A particle of mass 10kg is placed on a rough plane inclined at angle of 30degree to the horizontal. What horizontal force is required to move up the plane



1
Expert's answer
2022-03-18T09:08:40-0400

"F=\frac{mg\sin\alpha}{\cos\alpha}=mg\tan\alpha=57~N."